Scientific Notation (6.98 * 10-8th power)/ (3.67 * 10-5th power)
Divide 2 numbers/Express answer using scientific notation, retain at least three decimal places
1
Expert's answer
2012-03-28T16:10:10-0400
(6.98 * 10-8th power)/ (3.67 * 10-5th power) =(6.98 * 10^8)/ (3.67 * 10^5)=&
6.98/
3.67*10^(8-5)=1.902*10^3